### v2

# Optimal Modified Feedback Strategies in LQ Games under Control Imperfections

Abstract
Game-theoretic approaches and Nash equilibrium have been widely applied across various engineering domains. However, practical challenges such as disturbances, delays, and actuator limitations can hinder the precise execution of Nash equilibrium strategies. This work investigates the impact of such implementation imperfections on game trajectories and players’ costs in the context of a two-player finite-horizon linear–quadratic (LQ) nonzero-sum game. Specifically, we analyze how small deviations by one player, measured or estimated at each stage, affect the state and cost function of the other player. To mitigate these effects, we propose an adjusted control policy that optimally compensates for the deviations under the stated information structure and can, under certain conditions, exploit them to improve performance. Rigorous mathematical analysis and proofs are provided, and the effectiveness of the proposed method is demonstrated through a representative numerical example.
I Introduction
Game theory is a powerful analytical framework for modeling rational decision-making in complex multi-agent systems, particularly when interactions are strategic and agents’ decisions influence each other [1, 2]. Such interactions often involve conflicting objectives, motivating the development of solution concepts to characterize equilibrium behavior. Among these, the Nash equilibrium is a well-established notion, defined as a set of strategies from which no participant can unilaterally deviate to achieve a better outcome [3].
Thanks to its solid theoretical foundation, Nash equilibrium has been widely applied in diverse engineering domains, including robotics [4], traffic management [5], and sensor networks [6], as well as in emerging areas such as energy management [7, 8], autonomous vehicle navigation [9, 10, 11], and human–robot interaction [12].
Despite its wide adoption, implementing Nash strategies in real-world engineering systems faces significant challenges due to uncertainties, disturbances, delays, and actuator imperfections [13, 14, 15]. Extensions such as the $\varepsilon$-equilibrium relax optimality by allowing bounded deviations from Nash strategies [16], enabling numerical approximations [17, 18] and supporting further studies on sensitivity and robustness [13, 14, 19, 20, 21].
Robustness of Nash equilibrium to uncertainty has been studied from various perspectives. For example, model uncertainties in differential games were addressed in [13, 14, 21], where [13] designed a finite-horizon feedback Nash strategy for Linear Quadratic (LQ) games, [14] extended the scope to infinite-horizon settings, and [21] relaxed the LQ structure via weaker convexity assumptions. Other works addressed different uncertainty sources: [19] proposed adaptive control to mitigate measurement noise, and [20] analyzed the effect of cost parameter variations in discrete-time games.
Nonetheless, the sensitivity of Nash equilibrium to player-specific control deviations—arising from implementation imperfections—remains largely unexplored. Such deviations, e.g., due to actuator errors or delays, are common in engineering systems and can directly degrade performance. Moreover, the discrete-time setting has received little attention in this context. This work analytically investigates how such deviations influence system trajectories and player costs.
To address these effects, we propose a modified feedback policy that both mitigates performance loss and exploits available deviation information. While prior adaptive multi-agent methods exist [22, 23], they do not explicitly target compensation for execution-time control deviations of a specific player, which is addressed in this work.
This paper investigates finite-horizon, discrete-time LQ games in which one player’s control inputs deviate during execution. We analyze the resulting impact on joint trajectories and on the other player’s cost. The main contributions are summarized below:
  • A sensitivity analysis for discrete-time games, deriving closed-form expressions that characterize how one player’s deviations propagate through the system dynamics and influence the other player’s cost.
  • An optimal modified feedback control policy that compensates for such deviations and ensures performance that is never worse—and often strictly better—than the original feedback Nash equilibrium.
  • A complete theoretical analysis, supported by a representative numerical example, that highlights the robustness of Nash equilibrium under execution imperfections.
The remainder of the paper is organized as follows. Section II presents the problem formulation and background. Section III develops the sensitivity analysis. Section IV introduces the modified feedback control policy. Section V demonstrates its effectiveness through a numerical example. Finally, Section VI concludes the paper.
Notation. $\mathbb{R}$ and $\mathbb{Z}$ denote the sets of real and integer numbers, respectively. For a matrix $A$, $A^{\top}$ denotes its transpose; for a square matrix $B$, $B\succ 0$ ($B\succeq 0$) means $B$ is positive definite (positive semi-definite), and $B^{-1}$ denotes its inverse. For $i\in N_{p}=\{1,2\}$, the subscript -i refers to $N_{p}\setminus\{i\}$, i.e., all indices except $i$. The superscript denotes the exact optimal solution.
II Preliminaries
We focus on finite-horizon linear quadratic (LQ) games [24] with discrete-time dynamics
$x_{k+1}=Ax_{k}+B_{1}u_{1,k}+B_{2}u_{2,k},$ (1)
for $k\in\mathbb{K}=\{0,1,\dots,N-1\}\subset\mathbb{Z}$, where:
  • $x_{k}=\begin{bmatrix}x_{1,k}^{\top}&x_{2,k}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{1}+n_{2}}$ is the joint state vector obtained by concatenating the state vectors $x_{i,k}\in\mathbb{R}^{n_{i}}$ of Player $i\in\{1,2\}$, where $n_{i}$ is the dimension of Player $i$’s state.
  • $u_{1,k}\in\mathbb{R}^{m_{1}}$ and $u_{2,k}\in\mathbb{R}^{m_{2}}$ are the control inputs of Player 1 and 2, respectively, where $m_{i}$ is the dimension of Player $i$’s control input.
  • $A\in\mathbb{R}^{(n_{1}+n_{2})\times(n_{1}+n_{2})}$ is the joint state transition matrix.
  • $B_{1}\in\mathbb{R}^{(n_{1}+n_{2})\times m_{1}}$ and $B_{2}\in\mathbb{R}^{(n_{1}+n_{2})\times m_{2}}$ are the input matrices for Player 1 and 2, respectively.
Suppose Player $i\in\{1,2\}$ minimizes its cost $J_{i}$ over an $N$-step horizon by choosing control inputs $u_{i,k}$ at each stage $k\in\mathbb{K}$. The strategy of Player $i$ is $U_{i}=\{u_{i,k}\}_{k=0}^{N-1}$, and $U_{-i}$ denotes its counterpart’s strategy. The quadratic cost function of Player $i$ is
$J_{i}=\sum_{k=0}^{N-1}\left(x_{k}^{\top}Q_{i}x_{k}+u_{i,k}^{\top}R_{i}u_{i,k}\right)+x_{N}^{\top}Q_{i,N}x_{N},$ (2)
where $Q_{i},Q_{i,N}\succeq 0$ are the stage and terminal weighting matrices, and $R_{i}\succ 0$ is the control weighting matrix.
We use Nash equilibrium to mean a strategy profile $U_{i}$, where no player can reduce its cost by a unilateral deviation, i.e. $J_{i}(x_{0},U_{i}^{\star},U_{-i}^{\star})\leq J_{i}(x_{0},U_{i},U_{-i}^{\star})$ [3].
Definition 1.
Consider the LQ game (1)–(2). A pair of feedback laws
$u_{i,k}=-K^{\star}_{i,k}x_{k},\qquad i\in\{1,2\},\ k\in\mathbb{K},$
constitutes a linear Feedback Nash Equilibrium (FNE) if, for each $i$, $u_{i,k}$ minimizes $J_{i}$ against the opponent’s feedback $u_{-i,k}=-K^{\star}_{-i,k}x_{k}$ over the horizon.
The FNE gains $\{K^{\star}_{i,k}\}$ are obtained by a backward Riccati recursion coupled across players [25]. Let $\tilde{P}_{i,N}=Q_{i,N}$ and define, for each $k\in\{0,\dots,N-1\}$ (computed backward),
$A^{-i}_{\mathrm{cl},k}\;:=\;A-\sum_{j\neq i}B_{j}K_{j,k}.$
Given $\{\tilde{P}_{i,k+1}\}$ and $\{K_{-i,k}\}$, the matrices $\tilde{P}_{i,k}$ satisfy
$\tilde{P}_{i,k}=Q_{i}$ $+(A^{-i}_{\mathrm{cl},k})^{\top}\tilde{P}_{i,k+1}A^{-i}_{\mathrm{cl},k}$ (3)
$-(A^{-i}_{\mathrm{cl},k})^{\top}\tilde{P}_{i,k+1}B_{i}\,\tilde{H}_{k}\,^{-1}\!B_{i}^{\top}\tilde{P}_{i,k+1}A^{-i}_{\mathrm{cl},k},$
where $\tilde{H}_{k}:=R_{i}+B_{i}^{\top}\tilde{P}_{i,k+1}B_{i}$. The corresponding FNE feedback gains are
$K^{\star}_{i,k}=\tilde{H}_{k}^{-1}B_{i}^{\top}\tilde{P}_{i,k+1}A^{-i}_{\mathrm{cl},k}.$ (4)
Under the standard assumptions $Q_{i},Q_{i,N}\succeq 0$, $R_{i}\succ 0$ and nonsingularity of the gain system in Definition 1 at each $k$, the backward sweep $k=N-1\to 0$ yields a unique FNE [24, Sec. 6].
We next recall standard sufficient conditions that ensure the existence and uniqueness of the Nash equilibrium in Definition 1, thereby guaranteeing well-posedness of the game.
Remark 1.
Consider the following sufficient conditions for the existence of a unique Nash equilibrium in a two-player finite-horizon nonzero-sum discrete-time LQ game (2):
  1. 1.
    The state weighting matrices satisfy $Q_{i},Q_{i,N}\succeq 0$ and the control weighting matrices satisfy $R_{i}\succ 0$ for $i\in\{1,2\}$.
  2. 2.
[…]
  3. 3.
    Define
    $M_{k}=\begin{bmatrix}I&F_{1}&F_{2}\\ P_{1,k}&-I&0\\ P_{2,k}&0&-I\end{bmatrix},\quad F_{i}=B_{i}R_{i}^{-1}B_{i}^{\top},$
    where $I$ is the $n\times n$ identity and $P_{i,k}$ denote the player Riccati matrices from the standard backward recursions; invertibility of $M_{k}$ ensures well-posed coupled updates.
Under conditions 1-3, a unique finite-horizon Nash equilibrium exists. We assume these conditions henceforth [26].
III First-Order Sensitivity of Player 1 Cost under Opponent Deviations
Under the memoryless state-feedback information structure, a finite-horizon FNE for a two-player game has the form
$u^{\star}_{i,k}=-K^{\star}_{i,k}x_{k},\qquad i\in\{1,2\},\ \ k=0,\dots,N-1,$ (5)
as Definition 1. Substituting (5) into the dynamics (1) yields the time-varying closed-loop matrix
$A_{\mathrm{cl},k}:=A-B_{1}K^{\star}_{1,k}-B_{2}K^{\star}_{2,k},$ (6)
which evolves the dynamics
$x_{k+1}^{\star}=A_{\mathrm{cl},k}x_{k}^{\star},\qquad x_{0}^{\star}\equiv x_{0}.$ (7)
In practice, Player 2’s implemented input may deviate from its FNE command, $u_{2,k}=u^{\star}_{2,k}+\Delta u_{2,k}$. In this section, we quantify the effect of small deviations $\{\Delta u_{2,k}\}$ on Player 1’s cost when Player 1 keeps its FNE law (5). We obtain (i) an exact expression for the induced state perturbation and (ii) a first-order expansion of $\Delta J_{1}$ with explicit coefficients that respect the time variation in (6).
Definition 2.
For integers $l\geq k\geq j$, define
$\Phi(k,j):=\prod_{i=j}^{k-1}A_{\text{cl},i}\;=\;A_{\text{cl},k-1}\cdots A_{\text{cl},j},$
with the convention $\Phi(j,j):=I$. Then,
$\Phi(k+1,j)=A_{\text{cl},k}\,\Phi(k,j),\qquad\Phi(k,\ell)\,\Phi(\ell,j)=\Phi(k,j).$
Proposition 1.
Let $\{K^{\star}_{1,k},K^{\star}_{2,k}\}_{k=0}^{N-1}$ be the FNE gains in (5). Suppose Player 2 deviates as $u_{2,k}=u^{\star}_{2,k}+\Delta u_{2,k}$ while Player 1 keeps $u_{1,k}=u^{\star}_{1,k}=-K^{\star}_{1,k}x_{k}$. Then, for $k=1,\dots,N$,
$\Delta x_{k}\;=\;\sum_{j=0}^{k-1}\Phi(k,j+1)\,B_{2}\,\Delta u_{2,j},$ (8)
and, with $S_{k}:=Q_{1}+K_{1,k}^{\star\top}R_{1}K^{\star}_{1,k}$,
$\Delta J_{1}\;=\;2\sum_{j=0}^{N-1}x_{0}^{\top}\Lambda_{j}\,B_{2}\,\Delta u_{2,j}\;+\;\mathcal{O}\!\big(\|\Delta u_{2}\|_{2}^{2}\big),$ (9)
where
$\Lambda_{j}\;=\;\sum_{k=j+1}^{N-1}\Phi(k,0)^{\top}$ $S_{k}\,\Phi(k,j+1)$ (10)
$+\;\Phi(N,0)^{\top}Q_{1,N}\,\Phi(N,j+1),$
for $j=0,\dots,N-1$. Here $\Phi(k,j)$ is as in Definition 2, $\Delta u_{2}:=[\Delta u_{2,0}^{\top}\,\cdots\,\Delta u_{2,N-1}^{\top}]^{\top}$, and the remainder is with respect to the Euclidean norm $\|\Delta u_{2}\|_{2}$.
Proof.
Under the FNE , the nominal trajectory is obtained via (7). With $u_{2,k}=u_{2,k}^{\star}+\Delta u_{2,k}$ and $u_{1,k}=u_{1,k}^{\star}$, the actual state satisfies
$x_{k+1}\;=\;A_{\text{cl},k}\,x_{k}+B_{2}\,\Delta u_{2,k}.$
Defining $\Delta x_{k}:=x_{k}-x_{k}^{\star}$ gives
$\Delta x_{k+1}\;=\;A_{\text{cl},k}\,\Delta x_{k}+B_{2}\,\Delta u_{2,k},\qquad\Delta x_{0}=0,$
whose iteration yields (8) using $\Phi(k+1,j)=A_{\text{cl},k}\,\Phi(k,j)$.
Next, linearizing (2) for $i=1$ around the FNE trajectory $(x_{k}^{\star},u_{1,k}^{\star})$, leads to
$\Delta J_{1}\;=2\sum_{k=0}^{N-1}\big(x_{k}^{\star\top}$ $Q_{1}\Delta x_{k}+u_{1,k}^{\star\top}R_{1}\Delta u_{1,k}\big)$
$\;+$ $2\,x_{N}^{\star\top}Q_{1,N}\Delta x_{N}\;+\;\mathcal{O}(\|\Delta x\|_{2}^{2}),$
where $\Delta\,J_{1}=J_{1}(x_{k},u_{1,k})-J_{1}(x_{k}^{\star},u_{1,k}^{\star})$.
Since $u_{1,k}=-K^{\star}_{1,k}x_{k}$, we have $\Delta u_{1,k}=-K^{\star}_{1,k}\Delta x_{k}$, hence
$\Delta J_{1}\;=2\sum_{k=0}^{N-1}x_{k}^{\star\top}S_{k}\,\Delta x_{k}\;+2x_{N}^{\star\top}$ $Q_{1,N}\Delta x_{N}$ (11)
$+\;\mathcal{O}(\|\Delta x\|_{2}^{2}),$
where $S_{k}:=Q_{1}+K_{1,k}^{\star\top}R_{1}K^{\star}_{1,k}$. Substituting (8) into (11), defining $\Lambda_{j}$ as in (10), and using $x_{k}^{\star}=\Phi(k,0)x_{0}$ gives (10). As $\Delta x$ is linear in $\Delta u_{2}$ by (8), the remainder is $\mathcal{O}(\|\Delta u_{2}\|_{2}^{2})$. This completes the proof. ∎
Proposition 1 shows the deviations of Player 2 linearly perturb Player 1’s cost. If Player 1 keeps the equilibrium-derived feedback without compensation, e.g. $u_{1,k}^{\mathrm{REF}}:=-K^{\star}_{1,k}x_{k}$ (referred to as Reference Feedback (REF)), its policy is generally sub-optimal under the realized dynamics. Motivated by this sensitivity, we next design a controller that models the deviation and attenuates its effect as it propagates through the closed loop, compensating for these deviations.
IV Control Policy Design with Disturbance Compensation
In practice, the control implementation is rarely exact. Differences between the commanded and implemented inputs typically arise from (i) actuator/servo dynamics and inner-loop tracking (first-order lags, rate/slew limits), (ii) computation/communication latency and sample-and-hold effects, and (iii) command filtering/smoothing or safety layers that reshape inputs. This creates a measured input disturbance $w_{k}\;:=\;B_{2}\big(u_{2,k}-u_{2,k}^{\star}\big)$ in exactly the channel exposed by our sensitivity result.
We, therefore, model $w_{k}$ with simple, identifiable dynamics. Then, we equip Player 1 with a Compensated Feedback (CF) law, which includes a feedforward term $L_{k}w_{k}$ to cancel the predictable component of the disturbance at the actuator, so the feedback term $K_{k}$ regulates the residuals instead of reacting after the state has already moved. We first treat a physically grounded actuator-lag model, then narrow it down to a specific case of slow-varying error dynamics.
IV-A Compensated Feedback under Lagged Execution Error Dynamics
Assume that Player 2’s actuator is a discrete-time first-order lag that tracks its desired (Nash) input $u^{\star}_{2,k}=-K^{\star}_{2,k}x_{k}$ and let $u_{2,k}$ denote the implemented command. Let the execution error be $e_{k}:=u_{2,k}-u^{\star}_{2,k}$. Then, the plant disturbance is $w_{k}=B_{2}e_{k}$.
Assumption 1.
At time $k$, Player 1 observes $(x_{k},w_{k})$ and Player 2 implements its FNE command $u^{\star}_{2,k}=-K^{\star}_{2,k}x_{k}$ through a discrete first-order lag:
$u_{2,k+1}\;=\;\alpha\,u_{2,k}\;+\;(1-\alpha)\,u^{\star}_{2,k},\qquad\alpha\in(0,1).$ (12)
The induced disturbance recursion is
$w_{k+1}$ $\;=\;\alpha\,w_{k}\;-\;B_{2}\big(u^{\star}_{2,k+1}-u^{\star}_{2,k}\big)$ (13)
$\;=\;\alpha\,w_{k}\;+\;B_{2}\!\big(K^{\star}_{2,k+1}x_{k+1}-K^{\star}_{2,k}x_{k}\big).$
It is noteworthy that at each stage $k$, Player 1 knows the current $x_{k}$ and the current $w_{k}$, but not future values.
Remark 2.
The discrete recursion (12) is the zero-order-hold (ZOH) sampled counterpart of the continuous-time actuator $\dot{u}_{2}(t)=\tfrac{1}{\tau}(u_{2}^{\star}(t)-u_{2}(t))$ with sampling period $\Delta t$. The pole maps as
$\alpha\;=\;e^{-\Delta t/\tau},$ (14)
so the time constant $\tau$ of the continuous model sets the per-sample settling factor $\alpha\in(0,1)$.
The lag model stated in Assumption 1 captures two ubiquitous effects: (i) if the target $u^{\star}_{2,k}$ is steady, the tracking error decays geometrically with factor $\alpha$; (ii) changes in $u^{\star}_{2,k}$ inject a transient error that then dies out. Such behavior is widely used as a realistic low-order approximation of implementation effects.
Fix Player 2 at its Nash feedback $u^{\star}_{2,k}=-K^{\star}_{2,k}x_{k}$ and define
$A_{\mathrm{eff},k}:=A-B_{2}K^{\star}_{2,k},\qquad x_{k+1}=A_{\mathrm{eff},k}x_{k}+B_{1}u_{1,k}+w_{k}.$ (15)
Theorem 1.
Consider the game (2) with dynamics (1). Let Player 2 obey its FNE policy (5) and define the effective matrices $A_{\mathrm{eff},k}$ via (15). Under Assumption 1, let the disturbance $w_{k}$ be generated by the lag recursion (13). Assume $Q_{1},Q_{1,N}\succeq 0$ and $R_{1}\succ 0$. Over the causal affine class
$u_{1,k}\;=\;-K_{k}x_{k}\;-\;L_{k}w_{k},\qquad k=0,\dots,N-1,$ (16)
there exists a unique optimal policy that minimizes (2), with gains
$\big[\,K_{k}\ \ L_{k}\,\big]\;=\;\big(R_{1}+\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}\big)^{-1}\,\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k},$ (17)
where $\{\bar{P}_{k}\}$ is the unique sequence generated by the finite-horizon Riccati recursion
$\bar{P}_{N}=\bar{Q}_{N},$ (18)
$\begin{aligned} \bar{P}_{k}&=\bar{Q}+\bar{A}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k}\\ &-\bar{A}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}\big(R_{1}+\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}\big)^{-1}\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k},\end{aligned}$
$\text{for }k=N-1,\dots,0,$
on the augmented system with matrices
$\bar{A}_{k}=\begin{bmatrix}A_{\mathrm{eff},k}&I\\[2.0pt] B_{2}K^{\star}_{2,k+1}A_{\mathrm{eff},k}-B_{2}K^{\star}_{2,k}&\alpha I+B_{2}K^{\star}_{2,k+1}\end{bmatrix}$, (19)
$\bar{B}_{k}=\begin{bmatrix}B_{1}\\[2.0pt] B_{2}K^{\star}_{2,k+1}B_{1}\end{bmatrix}$,
where
$\bar{Q}=\begin{bmatrix}Q_{1}&0\\ 0&0\end{bmatrix},\quad\bar{Q}_{N}=\begin{bmatrix}Q_{1,N}&0\\ 0&0\end{bmatrix}$. (20)
Moreover, for any causal $\{\tilde{K}_{k},\tilde{L}_{k}\}$, the incurred cost satisfies the nonnegative gap identity
$\sum_{k=0}^{N-1}\!\big(\tilde{u}_{1,k}-u_{1,k}^{\star}\big)$ ${}^{\top}\big(R_{1}+\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}\big)\big(\tilde{u}_{1,k}-u_{1,k}^{\star}\big)$ (21)
$\,=\,J_{1}(\{\tilde{K}_{k},\tilde{L}_{k}\})-J_{1}(\{K_{k},L_{k}\})\ \geq 0,$
where $u_{1,k}^{\star}:=-K_{k}x_{k}-L_{k}w_{k}$ and $\tilde{u}_{1,k}:=-\tilde{K}_{k}x_{k}-\tilde{L}_{k}w_{k}$.
Proof.
Inspired by the Disturbance-Based State Compensation (DBSC) approach [27], which is commonly used in traditional state-feedback control design, and given Assumption 1, we define the augmented state as $z_{k}:=[x_{k}^{\top}\ w_{k}^{\top}]^{\top}$.
Then the game problem can be rewritten in augmented form
$\min_{\{u_{1,k}\}_{k=0}^{N-1}}z_{N}^{\top}\bar{Q}_{N}z_{N}+\sum_{k=0}^{N-1}\!\big(z_{k}^{\top}\bar{Q}z_{k}+u_{1,k}^{\top}R_{1}u_{1,k}\big),$ (22)
$\text{subject to:}\quad z_{k+1}=\bar{A}_{k}z_{k}+\bar{B}_{k}u_{1,k},\quad k=0,\dots,N-1,$ (23)
with $\bar{A}_{k},\bar{B}_{k},\bar{Q},\bar{Q}_{N}$ given in (19).
Define the value function of the augmented system
$V_{k}(z):=\min_{\{u_{1,\ell}\}_{\ell=k}^{N-1}}\Big\{z_{N}^{\top}\bar{Q}_{N}z_{N}+\sum_{\ell=k}^{N-1}\!(z_{\ell}^{\top}\bar{Q}z_{\ell}+u_{1,\ell}^{\top}R_{1}u_{1,\ell})\Big\}$ (24)
subject to the augmented dynamic constraint of (23).
Based on the Bellman principle of optimality at time $k$, we can reformulate $V_{k}(z)$ as
$V_{k}(z_{k})=\min_{u_{1,k}}\Big\{z_{k}^{\top}\bar{Q}z_{k}+u_{1,k}^{\top}R_{1}u_{1,k}+V_{k+1}\big(z_{k+1}\big)\Big\},$ (25)
Backward induction yields $V_{k}(z_{k})=z_{k}^{\top}\bar{P}_{k}z_{k}+\bar{c}_{k}$ with $\bar{P}_{k}$ given by (18). Substituting the augmented dynamics (23) in (25) and collecting terms in $u_{1,k}$ yields the strictly convex quadratic in $u_{1,k}$:
$\mathcal{Q}_{k}(z_{k},u_{1,k})=u_{1,k}^{\top}$ $\mathcal{H}_{k}\,u_{1,k}+2\,u_{1,k}^{\top}\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k}\,z_{k}$ (26)
$+z_{k}^{\top}\!\big(\bar{Q}+\bar{A}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k}\big)z_{k}+\bar{c}_{k+1},$
where $\mathcal{H}_{k}:=R_{1}+\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}$ which is positive definite as $R_{1}\succ 0$ and $\bar{P}_{k+1}\succeq 0$.
Hence, the unique minimizer is
$u_{1,k}\;=\;-\mathcal{H}_{k}^{-1}\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k}\,z_{k},=:-\big[\,K_{k}\ \ L_{k}\,\big]z_{k}$ (27)
which, upon partitioning $z_{k}=[x_{k}^{\top}\;w_{k}^{\top}]^{\top}$, gives (17).
Completing the square yields (18). Summing the stagewise completion terms from $k=0$ to N1N-1 telescopes the value functions and proves (21). Uniqueness follows the strict convexity ($\mathcal{H}_{k}\succ 0$). ∎
By (21) in Theorem 1, the gains $\{K_{k},L_{k}\}$ are optimal over the causal affine class (16), which includes the REF baseline $u_{1,k}^{\text{REF}}$ as the special case $L_{k}\equiv 0$. Consequently, execution-time deviations by Player 2 create an opportunity for Player 1 to outperform its own equilibrium-derived feedback (REF) by adopting the compensated law. This does not contradict the definition of a Nash equilibrium: the FNE is defined for the nominal game without implementation deviations, whereas here Player 2’s realized input differs from its Nash command. The approach is practically relevant because such deviations routinely arise from disturbances, physical limits, and execution imperfections.
The next result quantifies the local benefit, giving a signed, second-order improvement in the deviation magnitude.
Corollary 2.
Let $J_{1}^{\mathrm{CF}}$ and $J_{1}^{\mathrm{REF}}$ be the costs under $u_{1,k}^{\mathrm{CF}}=-K_{k}x_{k}-L_{k}w_{k}$ and $u_{1,k}^{\mathrm{REF}}=-K^{\star}_{1,k}x_{k}$, respectively, where $\{\bar{P}_{k}\}$ and $K_{k}\ \ L_{k}\,]$ are as in Theorem 1. Define the selector $E_{w}:=[\,0\ \ I\,]$ so that $w_{k}=E_{w}z_{k}$ for $z_{k}:=[x_{k}^{\top}\ w_{k}^{\top}]^{\top}$. For sufficiently small deviations $\{\Delta u_{2,k}\}$ where $w_{k}=\mathcal{O}(\|\Delta u_{2}\|_{2})$,
$J_{1}^{\mathrm{CF}}-J_{1}^{\mathrm{REF}}\;=\;-\,\sum_{k=0}^{N-1}\Big\|\mathcal{H}_{k}^{-\tfrac{1}{2}}\,\bar{B}_{k}^{\top}$ $\bar{P}_{k+1}E_{w}\,w_{k}\Big\|_{2}^{\,2}$ (28)
$\;+\;\mathcal{O}\!\big(\|\Delta u_{2}\|_{2}^{3}\big)\;\leq 0,$
where $\mathcal{H}_{k}:=R_{1}+\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}\succ 0$. Thus, CF never performs worse than REF to second order in the deviation magnitude and strictly improves unless $w_{k}\equiv 0$.
Proof.
By Theorem 1, the stagewise completion of squares gives, for any candidate $\tilde{u}_{1,k}$, $\mathcal{Q}_{k}(z_{k},\tilde{u}_{1,k})-\mathcal{Q}_{k}(z_{k},u_{1,k}^{\star})=\|\mathcal{H}_{k}^{1/2}(\tilde{u}_{1,k}-u_{1,k}^{\star})\|_{2}^{2}$. Evaluating this at the REF choice $\tilde{u}_{1,k}=-K_{1,k}^{\star}x_{k}$ yields $\tilde{u}_{1,k}-u_{1,k}^{\star}=L_{k}w_{k}$ plus higher-order state terms induced by $\Delta u_{2}$, which are $\mathcal{O}(\|\Delta u_{2}\|_{2}^{2})$. Summing over $k$ gives (28) with the remainder $\mathcal{O}(\|\Delta u_{2}\|_{2}^{3})$ and completes this proof. ∎
IV-B Specialization: Slowly-Varying Error Dynamics
When the FNE command $u_{2,k}^{\star}$ varies slowly, the drive term becomes negligible and $w_{k+1}=\alpha w_{k}$. The CF gains then admit closed forms in terms of the block Riccati sequence.
Corollary 3.
Assume the lag model (13) with $0<\alpha<1$ and that Player 2’s Nash command varies slowly so that
$u^{\star}_{2,k+1}\approx u^{\star}_{2,k}$ (29)
hence the drive is negligible and $w_{k+1}=\alpha\,w_{k}$. Then the optimal gains $K_{k}$ and $L_{k}$ defined in Theorem 1 will be obtained as
$K_{k}\;=\;H_{k}^{-1}B_{1}^{\top}P_{k+1}^{xx}A_{\mathrm{eff},k},$ (30)
$L_{k}\;=\;H_{k}^{-1}B_{1}^{\top}\big(P_{k+1}^{xx}+\alpha\,P_{k+1}^{xw}\big),$
where $H_{k}\;:=\;R_{1}+B_{1}^{\top}P_{k+1}^{xx}B_{1}\ \succ 0$; $P_{k}^{xx}$ satisfies the standard finite-horizon Riccati recursion
$P_{N}^{xx}=Q_{1,N},$ (31)
$\begin{aligned} P_{k}^{xx}=Q_{1}+&A_{\mathrm{eff},k}^{\top}P_{k+1}^{xx}A_{\mathrm{eff},k}\\ &-A_{\mathrm{eff},k}^{\top}P_{k+1}^{xx}B_{1}H_{k}^{-1}B_{1}^{\top}P_{k+1}^{xx}A_{\mathrm{eff},k},\end{aligned}$
and the cross block $P_{k}^{xw}$ evolves linearly as
$P_{N}^{xw}=0,$ (32)
$\begin{aligned} P_{k}^{xw}=\;&A_{\mathrm{eff},k}^{\top}P_{k+1}^{xx}+\alpha\,A_{\mathrm{eff},k}^{\top}P_{k+1}^{xw}\\ &-A_{\mathrm{eff},k}^{\top}P_{k+1}^{xx}B_{1}H_{k}^{-1}B_{1}^{\top}\big(P_{k+1}^{xx}+\alpha\,P_{k+1}^{xw}\big).\end{aligned}$
Proof.
The assumption (29) implies $w_{k+1}=\alpha w_{k}$. In augmented coordinates $z_{k}=[x_{k}^{\top}\;w_{k}^{\top}]^{\top}$ this gives
$\bar{A}_{k}=\begin{bmatrix}A_{\mathrm{eff},k}&I\\ 0&\alpha I\end{bmatrix},\qquad\bar{B}_{k}=\begin{bmatrix}B_{1}\\ 0\end{bmatrix}$.
By Theorem 1, the optimal control (27) can be simplified by partition $\bar{P}_{k+1}=\begin{bmatrix}P_{k+1}^{xx}&P_{k+1}^{xw}\\ P_{k+1}^{wx}&P_{k+1}^{ww}\end{bmatrix}$ and computing the terms
$\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{B}_{k}\;=\;B_{1}^{\top}P_{k+1}^{xx}B_{1},$
$\bar{B}_{k}^{\top}\bar{P}_{k+1}\bar{A}_{k}=B_{1}^{\top}\!\big[\,P_{k+1}^{xx}A_{\mathrm{eff},k}\ \ \ P_{k+1}^{xx}+\alpha P_{k+1}^{xw}\,\big].$
Defining $H_{k}:=R_{1}+B_{1}^{\top}P_{k+1}^{xx}B_{1}\succ 0$,
$u_{1,k}^{\star}$ $=-\,H_{k}^{-1}\,B_{1}^{\top}\!\big[\,P_{k+1}^{xx}A_{\mathrm{eff},k}\ \ \ P_{k+1}^{xx}+\alpha P_{k+1}^{xw}\,\big]\begin{bmatrix}x_{k}\\ w_{k}\end{bmatrix}$ (33)
$\;\begin{aligned} =&-\,\underbrace{H_{k}^{-1}B_{1}^{\top}P_{k+1}^{xx}A_{\mathrm{eff},k}}_{\displaystyle K_{k}}\,x_{k}\\ &\;-\;\underbrace{H_{k}^{-1}B_{1}^{\top}\big(P_{k+1}^{xx}+\alpha P_{k+1}^{xw}\big)}_{\displaystyle L_{k}}\,w_{k},\end{aligned}$
which yields (30). Taking the $(1,1)$ and $(1,2)$ blocks of the Riccati recursion (18) with the above $\bar{A}_{k},\bar{B}_{k}$ gives the stated recursions (31)–(32) and $P_{N}^{xw}=0$ follows from $\bar{Q}_{N}$ being block-diagonal. ∎
Comparing the Riccati recursion $P^{xx}_{k}$ in (31), and the gain formulation $K_{k}$ in (30) with the ones in Definition 1 reveals that the state gain $K_{k}$ obtained under the assumption of Corollary 3 is equal to that of the FNE.
It is worth noting that (31) coincides with the standard finite-horizon Riccati recursion for the disturbance-free system. In particular, $K_{k}$ matches the usual state-feedback gain derived from $P_{k}^{xx}$, while $L_{k}$ adds a causal feedforward term shaped by both $P_{k+1}^{xx}$ and the cross block $P_{k+1}^{xw}$ scaled by $\alpha$. The resulting law $u_{1,k}^{\star}=-K_{k}x_{k}-L_{k}w_{k}$ thus has the familiar structure of disturbance-accommodation or preview control: the controller reacts not only to the state but also to the measured disturbance. Under the lag generator (13), the augmented Riccati (18) introduces off-diagonal couplings between $x$ and $w$; these generate a preview component in $L_{k}$ that anticipates how current deviations propagate through the lag into future stages. Hence (17) recovers the classical disturbance-accommodation principle and extends it to physically grounded, time-varying deviation dynamics.
Remark 3.
To sanity-check the design, we examine the two extreme points of the lag parameter $\alpha$ in the implemented-command model (12). When $\alpha\to 0$, the actuator tracks its Nash command essentially instantaneously, so $u_{2,k+1}\approx u^{\star}_{2,k}$ and the disturbance recursion reduces to $w_{k+1}=-B_{2}(u^{\star}_{2,k+1}-u^{\star}_{2,k})$. In steady segments the disturbance vanishes, and the compensated feedback coincides with the classical disturbance-accommodation (DBSC) law that uses $w_{k}$ as an immediate feedforward signal. At the other extreme, when $\alpha\to 1$ and the drive term is negligible, the disturbance behaves like a nearly constant bias with $w_{k+1}\approx w_{k}$, so the compensation acts as preview cancellation of a persistent input. Both cases are contained as limits of the proposed lag model and show that the compensated law weakly dominates any $K$-only feedback whenever a measurable deviation is present.
Remark 4.
Under the CF policy for Player 1 (16) and the feedback policy of Player 2 (5), the closed-loop state update can be written as
$x_{k+1}=\big(A-B_{1}K_{1,k}-B_{2}K^{\star}_{2,k}\big)\,x_{k}-B_{1}L_{k}w_{k}.$
Hence, even for $w_{k}\equiv 0$, the homogeneous (zero-disturbance) dynamics are
$x_{k+1}=\big({A-B_{1}K_{1,k}-B_{2}K^{\star}_{2,k}}\big)\,x_{k},$
which generally differ from the FNE dynamics $x_{k+1}=\big(A-B_{1}K^{\star}_{1,k}-B_{2}K^{\star}_{2,k}\big)\,x_{k}$. Stability and performance must therefore be assessed for the Linear Time Invariant (LTV) matrix sequence produced by the augmented Riccati recursion.
It is noteworthy that the CF policy is not a Nash equilibrium for the modified interaction, since Player 2 is not best-responding to CF. Joint redesign of both players under implementation dynamics is more involved and is left for future work.
The next section presents a numerical example illustrating the practical utility of Theorems 1.
V Numerical Example: Two Carts Coupled by a Spring–Damper
We consider a two-player, finite-horizon, discrete-time LQ dynamic game with feedback Nash equilibrium (FNE) strategies. Two equal-mass carts are connected by a spring–damper; Player 1 and Player 2 actuate Cart 1 and Cart 2, respectively. Each player regulates its cart to the origin while penalizing control effort. We then impose a realistic execution-lag deviation at Player 2 and show that Player 1 can employ the CF from Theorem 1, computed via an augmented Riccati recursion, to counteract that structured disturbance while Player 2 remains at FNE.
Let $x:=\begin{bmatrix}p_{1}&v_{1}&p_{2}&v_{2}\end{bmatrix}^{\top}$ in which $p_{1}$ and $p_{2}$ are the positions and $v_{1}$ and $v_{2}$ are the velocities of Player 1 and 2, respectively. The continuous-time system dynamics are $\dot{x}=A_{c}x+B_{1,c}u_{1}+B_{2,c}u_{2}$ with the matrices:
$A_{c}=\begin{bmatrix}0&1&0&0\\[2.0pt] -\frac{k}{m}&-\frac{c}{m}&\frac{k}{m}&\frac{c}{m}\\[2.0pt] 0&0&0&1\\[2.0pt] \frac{k}{m}&\frac{c}{m}&-\frac{k}{m}&-\frac{c}{m}\end{bmatrix}$,
$B_{1,c}=\begin{bmatrix}0\\ \frac{1}{m}\\ 0\\ 0\end{bmatrix},\quad B_{2,c}=\begin{bmatrix}0\\ 0\\ 0\\ \frac{1}{m}\end{bmatrix}$.
We discretize the dynamics via forward Euler at sampling $\Delta T$. Each player minimizes the quadratic cost in (2) over a horizon $N$ with the weights $Q_{1}=\mathrm{diag}(10,\,2,\,0,\,0)$, $Q_{2}=\mathrm{diag}(0,\,0,\,10,\,2)$, and $R_{1}=R_{2}=[\,0.05\,]$.
For error identification purposes, let $u_{2,k}^{\star}=-K_{2,k}^{\star}x_{k}$ denote Player 2’s intended (FNE) command, and let $u_{2,k}$ be the applied input after a first-order discrete lag (12) with the $\alpha$ defined in Remark 2. The parameters used are listed in Table I.
Table I: Parameters used in the numerical simulations.
Model parameters
$m$ $1.0$ Cart mass [kg]
$k$ $0.12$ Spring constant [N/m]
[…]
$\Delta T$ $0.015$ Sampling period [s]
NN $40$ Horizon length [steps]
Error/lag parameters
$\tau$ $0.8$ Lag time constant [s]
$\alpha$ $e^{-\Delta T/\tau}\approx 0.9814$ Discrete lag factor [–]
Initial conditions
$x_{0}$ $\begin{bmatrix}0.5&0&-0.5&0\end{bmatrix}^{\top}$ Initial state [m, m/s, m, m/s]
Test Cases
We compare three implementations over the same horizon and initial condition. Case I (FNE): both players apply their finite-horizon feedback Nash laws computed from the coupled Riccati recursion; no actuator lag is present, and Player 2’s intended command is applied directly. Case II (REF): Player 2’s intended command is subjected to a first-order discrete lag with factor $\alpha$, but Player 1 continues to use its nominal feedback policy; this isolates the performance loss due to the execution lag. Case III (CF): the same actuator lag is present on Player 2, while Player 1 designs a compensated feedback by solving a finite-horizon Riccati recursion on the exact augmented realization of the lagged plant, yielding time-varying gains that act on the state and on the lag-induced mismatch signal.
Figure 1: Schematic of the two-cart setup and qualitative final positions under the three cases. Case I (FNE) shows near-symmetric convergence near the origin. Case II (REF) illustrates how Player 2’s actuator lag degrades both players’ positions relative to nominal. Case III (CF) shows that the compensated policy enables Player 1 to mitigate the error and approach its nominal outcome, while Player 2 remains misaligned due to its uncompensated lag.
Figure 2 compares the state and input trajectories under the three scenarios. In the nominal FNE (Case I, solid lines), both players regulate their carts symmetrically to the origin with smooth, coordinated actions. With a first-order execution lag at Player 2 but no compensation (Case II, dashed), the applied input $u_{2}$ deviates from the intended FNE signal (step-like due to the zero-order hold), producing a sluggish response in $(p_{2},v_{2})$ and, through the spring–damper coupling, a degraded transient in $(p_{1},v_{1})$. This mismatch yields a significant cost increase for both players (Table II).
Under the compensated feedback (Case III, dotted), Player 1 modifies its input relative to the nominal FNE action to proactively counteract the actuation shortfall of Player 2. The resulting trajectories stay appreciably closer to the FNE benchmark—especially for Player 1—confirming that CF recovers a substantial fraction of the lag-induced performance loss. Player 2’s cost, however, rises further because it neither alters its strategy nor removes the actuator lag, highlighting the asymmetric benefit in a general-sum game.
Numerically (Table II), Player 1’s cost increases by $51.36\%$ when lag is present without compensation (Case II vs. Case I). With CF (Case III), this penalty is reduced: $J_{1}$ drops from $492.40$ to $439.67$, a $10.71\%$ improvement relative to the uncompensated case, recovering about $31.56\%$ of the loss from the nominal baseline.
Figure 2: State trajectories $(p_{1},v_{1},p_{2},v_{2})$ and applied inputs under the three cases. The compensated feedback (CF) attenuates the effect of Player 2’s actuator lag by shaping a mismatch-aware control for Player 1.
Table II: Finite-horizon costs under the three cases ($\tau=0.8$ s, $\alpha\approx 0.9814$). CF (Case III) reduces Player 1’s lag penalty relative to Case II.
Case I: FNE Case II: Lag, no comp. Case III: CF
$J_{1}$ 325.31 492.40 439.67
$J_{2}$ 330.45 492.11 556.40
VI Conclusion
In this paper, we analyzed the impact of control implementation errors in finite-horizon, two-player, nonzero-sum LQ games. Our analysis focused on how deviations from the Nash equilibrium influence system stability and performance. We quantified the deviation of the system trajectory from the Nash trajectory and assessed the increase in the cost function when one player fails to follow the exact equilibrium strategy. Our analysis demonstrated that, when a Nash equilibrium exists, the system trajectory remains close to the Nash trajectory for small deviations from the equilibrium strategy.
Additionally, we proposed an adjusted control policy to mitigate the adverse effects of deviations by the other player from the Nash policy. This adjustment, grounded in the principle of dynamic programming, offers superior performance compared to strictly adhering to the pure Nash strategy. By optimally modifying the player’s response, the proposed policy minimizes the negative impact of the other player’s deviations and capitalizes on these deviations to achieve enhanced outcomes. The effectiveness of the proposed approach was validated through a known example. These findings highlight the potential for improved performance in real-world applications where executing the exact Nash strategy is impractical due to practical challenges such as actuator limitations or external disturbances.
Future Research Directions
A promising avenue for future research is the development of online estimation techniques combined with adaptive control laws that can anticipate deviations of the other player from the Nash strategy and dynamically adjust to these errors in real time. Furthermore, examining the effects of more complex deviation patterns—such as time-varying, correlated, or structured errors—could offer deeper insights into the robustness and adaptability of game-theoretic control strategies.
VII Acknowledgment
The authors acknowledge the use of OpenAI’s ChatGPT for structuring content, refining language, and improving clarity. All technical contributions, derivations, and original insights remain the authors’ own work.